1) A student says that 13 * 4 + 6 is the same as 13 * 10. How do you respond?
Found 2 solutions by Mathtut, MathTherapy:
Answer by Mathtut(3670)

I would say the are absolutely wrong.
:
according to the correct order of operations
:
13*4+6 would be 13 times 4 and then +6--->13(4)+6=58--->multiplication takes place before addition and subtraction in order from left to right
:
13*10=130

Answer by MathTherapy(10549)

1) A student says that 13 * 4 + 6 is the same as 13 * 10. How do you respond?
This student is very wrong and needs to spend time studying "Order of Operations," which has the acronym PEMDAS. With that said, the student needs to do the multiplication part of the problem before adding the result to 6, or
13 * 4 + 6 = (13 * 4) + 6 = 52 + 6 = 58
Done the wrong way, the result will be 130 (13 * 10). This is very incorrect.